# SAM 2++: Tracking Anything at Any Granularity

### v3

@@ 9 Data Details
The key features of this dataset are as follows: (1) High Resolution: The dataset consists of high-resolution videos, ensuring that fine details are preserved and enabling more accurate analysis. (2) Diversity: It encompasses a wide variety of scenes, sources, and tracked object categories, providing a rich and representative sample of real-world scenarios. (3) Complex and Challenging Cases: The dataset includes numerous complex situations, such as occlusion, motion blur, and other challenging visual conditions, which test the robustness and generalization ability of tracking algorithms. (4) Comprehensive Annotations: the dataset contains annotations at multiple granularities, including segmentation masks, bounding boxes, and key points.
9.1 Data Requirements
Videos Requirements. The selected videos must satisfy the following criteria:
  • No camera cuts or scene transitions are present throughout the video.
  • Visuals are clear, and the boundaries of the target can be accurately identified.
  • The duration is between 10 and 40 seconds (excluding static images).
  • Each video must contain at least one target object that meets the outlined below.
Target Object Requirements. Each video must include at least one object designated as the tracking target, which must fulfill the following basic criteria:
  • The target has clearly distinguishable boundaries from other objects in the scene.
  • Eligible targets include the full body or parts of a human (e.g., face, facial features, limbs, hands, feet, etc.) or an animal (full body or parts).
  • The target must appear in the first frame of the video and be clearly identifiable.
  • At least one key point on the target must be visible and locatable for most of the video, allowing brief occlusions or exits.
  • The target should be in motion (either actively or passively) for most of the video.
  • To ensure the dataset emphasizes challenging tracking scenarios, the target must also meet at least one of the following additional difficulty criteria:
    • Rapid movement of the target itself or due to camera motion.
    • High similarity to other objects.
    • Occlusion or brief disappearance and reappearance.
    • Deformation (e.g., shape or structure changes) or notable changes in size, orientation, or viewpoint (e.g., approaching or turning).
    • The target is small relative to the frame, but not excessively tiny.
Target Point Requirements. We further pick at least one point on the chosen target object. These points need to meet the following conditions:
  • The point could be the center point, the corner point, or semantically meaningful points such as human eyes, hands, or head.
  • The key point must be present in the first frame.
  • If the key point is occluded or disappeared, it should be labeled as ”occluded.”
  • For spherical objects, the key point should be placed near the center.
Many current point tracking datasets use arbitrary points as annotation targets. However, we chose to focus on keypoints as the target for both data annotation and model optimization for the following reasons, primarily based on two considerations: 1) Practical Application Perspective: Downstream tasks like 3D reconstruction and SLAM, require tracking key points in most cases. Key points offer stronger distinguishing and descriptive capabilities, and typically only a small number of high-quality key points are sufficient for other tasks, eliminating the need to track any-point. 2) Annotation Cost Efficiency: Annotating any-point incurs prohibitively high costs. Unlike RoboTAP (which relies on optical-flow-based trajectory interpolation and is limited to lab scenes) or Kubric/RGB-Stacking (which generates point annotations via rendering, lacking real-world diversity), our dataset sources videos from indoor, outdoor, and wild environments from real-world. To ensure annotation accuracy, the target points were selected by annotators and manually labeled frame-by-frame. Due to the unbearable time and human resources required for any-point annotation, keypoint annotation is a better choice to balance dataset utility and feasibility. Similarly, real-scene datasets like DAVIS and Kinetics annotate most salient objects.
Table 9: Performance Comparison of Automatic Visible Annotation.
@@ 9.1 Data Requirements
Table 10: Performance Comparison of Automatic Annotation in Different Annotation Methods.
(a) mask automatic annotation
#1 94.4 91.7 97.1 94.6 91.9 97.3
#2 95.0 92.4 97.6 95.0 92.4 97.6
#3 94.5 91.8 97.2 94.7 92.0 97.4
@@
(b) box automatic annotation
Type + 0th Back. AUCval $P_{Norm}^{val}$ $P^{val}$ AUCtest $P_{Norm}^{test}$ $P^{test}$
@@
#2 83.7 93.9 91.3 84.5 94.7 91.0
#3 83.9 93.8 91.3 84.7 94.5 91.0
#4 83.6 93.6 91.1 84.4 94.4 90.9
(c) point automatic annotation
+ 0th Back. Vis. OAval AJval OAtest AJtest
#1 - 89.0 61.7 88.7 61.3
[…]
#3 OR 89.3 58.1 89.6 57.3
#4 AND 88.2 62.1 87.7 61.1
#5 OR 89.4 58.8 89.6 57.7
#6 AND 88.4 62.6 87.8 61.4
Figure 7: Example videos from the Tracking-Any-Granularity dataset with annotation at various granularities. Each annotation has a unique color. Better viewing with zoom and color.
9.2 Annotation Pipeline
We designed a coarse-to-fine annotation pipeline to ensure high-quality multi-granularity annotations, which consists of the following four steps.
1) Video Selection. We downloaded a large number of videos from YouTube and instructed the annotators to select videos and objects that meet the above requirements.
2) Coarse Annotation. Annotators mark key points and tight bounding boxes on target objects.
3) Fine Annotation. To reduce annotator workload and improve efficiency, we use SAM [kirillov2023seganysam1] to generate rough masks based on the coarse annotations (points and boxes). Then, annotators refine these masks with the following requirements:
  • Only annotate the visible parts of the present object.
  • In cases of motion blur, infer the approximate position based on the previous frame to maintain temporal consistency. Masks in adjacent frames should not differ drastically.
  • Ignore transparent or semi-transparent watermarks and subtitles when creating masks; masks can directly cover these elements.
  • Exclude opaque overlays (such as logos or captions) from the mask.
  • For containers holding other objects, do not include the contained objects in the mask.
  • The mask should tightly fit the object, neither exceeding nor falling short of its boundaries.
  • Ensure that mask edges are smooth and avoid excessive roughness.
  • Fill in small internal holes, but preserve natural gaps (such as hollowed-out structures) or occlusions caused by other objects.
  • If the initial SAM-generated mask is of very poor quality, annotators may clear it entirely and use color tolerance-based selection to manually annotate the object from scratch.
4) Final Completion. Experts perform a final review to thoroughly assess the accuracy and consistency of all three types of annotations, ensuring that the labeling meets the required standards and that any discrepancies are identified and corrected.
9.3 Data engine
To increase the size of the dataset while reducing the workload, we adopted a selective annotation strategy in the second and third phases. Instead of manually labeling every video frame, annotators labeled only a subset of frames at varying intervals. After training the model on both public datasets and the fully labeled data from earlier phases, we leveraged the model to automatically annotate the remaining frames. Specifically, each video was divided into multiple clips, with annotators manually labeling the first and last frames of each clip. The annotation of the first frame in each clip served as the initial target state, enabling the model to infer the target state in the intermediate frames.
To further enhance annotation quality, we introduced two optional refinement methods: (1) performing backward tracking and fusing the results with those from forward tracking, and (2) since the target may be absent in some annotated frames, using the first frame of the entire video, which is guaranteed to contain the target, as an additional reference state alongside the first frame of each clip. We evaluated these enhancement methods on the Phase 1 validation and test set to determine the inference setting for each tracking task, as shown in Table 10. Specifically, (a), (b), and (c) represent the evaluation outcomes for the VOS, SOT, and PT tasks under various settings, respectively, while Table 9 shows the evaluation results for object existence prediction. By comparing the results across different settings, we select the configuration highlighted in the gray row as the inference setting for each task.
Table 11: Statistical analysis of video data from our dataset.
@@ 9.3 Data engine
Table 12: Attribute analysis of video data from our dataset.
Attribute. Definition Num.
BC. Background Clutter. The appearances of background and target object are similar. 1284
CS. Camera-Shake. Footage displays non-negligible vibrations. 429
DB. Dynamic Background. Large movement of background areas or other objects. 1523
DEF. Deformation. Target appearance deform complexly. 3580
EA. Edge Ambiguity. Unreliable edge detection, such as thorny sea urchins and rolling waves. 339
FO. Full Occlusion. Object becomes fully occluded, accompanied by Partially Occlusion in most cases. 253
FOV. Fully Out of View. The object is fully clipped by the image boundaries. 203
IV. Illumination Variation. when illumination in object region heavily varies. 232
MB. Motion Blur. Boundaries of target object is blurred because of camera or object fast motion. 4332
OP. Object Part. The object is a part of the whole. 761
PO. Partially Occlusion. The object becomes partially occluded. 2844
POV. Partially out of view. The object is partially clipped by the image boundaries. 1498
ROT. Rotation. The object rotates. 2279
$R$. Reflected and Refraction. The object undergoes reflection or refraction. 267
SA. Similar Appearance. There are multiple different objects that are similar to the target object. 1546
SC. Shape Complexity. Boundaries of target object is complex. 123
TO. Transparent Object. The object is transparent 87
VC. Viewpoint Change. The camera viewpoint changes. 1304
@@ 9.4 Tracking-Any-Granularity Dataset
Our dataset comprises 6,000 videos, each annotated with three types of labels: masks, boxes, and points. Fig. 7 shows some videos with various annotations.
Statistics and Attribute. The resolution of the majority of the videos is $1280\times 720$, with 398 exceptions. The duration of the videos ranges from $5.3$ seconds to $110.9$ seconds, and the frame count varies from $80$ frames to $3,317$ frames. In total, the dataset comprises $2.2$ million frames, amounting to a cumulative duration of $1,338.7$ minutes. More detailed statistics are shown in Table 11. We label each sequence with 18 attributes that represent various video challenges, as shown in Table 12.
10 Additional Experiments
10.1 Performance Comparison
Evaluation metrics. In video object segmentation task, we use standard metrics [vos-benchmark] in most benchmarks: Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$. In the YouTubeVOS benchmark, $\mathcal{J}$ and $\mathcal{F}$ are computed for ”seen” and ”unseen” categories separately. $\mathcal{G}$ is the averaged $\mathcal{J\&F}$ for both seen and unseen classes. In LVOS benchmark, the first densely annotated long-term VOS dataset with high-quality annotations, it introduces the standard deviation $\mathcal{V}$ of the average score of $\mathcal{J}$ and $\mathcal{F}$ to assess the temporal stability of VOS models. In VOST benchmark, which focuses on segmenting objects as they undergo complex transformations, it additionally reports $\mathcal{J}_{tr}$ for the last 25% of the frames in a sequence to show the robustness after the transformation has been mostly completed. The BURST benchmark is evaluated with Higher Order Tracking Accuracy (HOTA) [hota] as a good balance between measuring frame-level detection and temporal association accuracy. In single object tracking task, we evaluate performance with Area Under the Curve (AUC), normalized precision (${P}_{Norm}$) and precision (P{P}) to measure the average accuracy of center, size and scale between the prediction and labeled groundtruth bounding boxes of all the frames for most benchmarks. For the GOT-10k benchmark, we choose the average overlap (AO) and success rate (SR) as indicators. The former AO denotes the average of overlaps between all groundtruth and estimated bounding boxes, while the SR measures the percentage of successfully tracked frames where the overlaps exceed a threshold (e.g., 0.5). In point tracking task, we report Occlusion Accuracy (OA) and Average Jaccard (AJ) for TAP-Vid [tapvid], Perception Test [patraucean2023perception], and our dataset. As for the BADJA benchmark, we adopt the Percentage of Correct Keypoint-Transfer (PCK-T). We measure these benchmarks in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries.
Figure 8: Examples from SAM 2++ results on video benchmarks at various granularities.
Figure 9: Comparison between our model and various SOTA methods on video tracking benchmarks at three granularities. Better viewing with zoom and color.
Table 13: Comparison with Unified Models.

### v4

@@ 5 Tracking-Any-Granularity Dataset
We developed a comprehensive dataset for training our unified model, termed Tracking-Any-Granularity (TAG), with annotations across three granularities: segmentation masks, bounding boxes, and key points. Our dataset contains 6,000 high-resolution videos, which preserve fine visual details and enable accurate tracking analysis. It further encompasses diverse scenes, video sources, and tracked object categories, providing a representative sample of real-world scenarios. Moreover, TAG includes complex and challenging cases, such as occlusion, motion blur, and other difficult visual conditions, to test the robustness and generalization ability of tracking algorithms. With comprehensive multi-granularity annotations, a three-phase data engine with model-in-the-loop annotation workflows, and strict multi-stage quality checks, we ensure large-scale, high-quality, and consistent annotations.
5.1 Target Selection
Before annotation, we define strict selection criteria for videos, target objects, and target points to ensure data quality and tracking difficulty. For videos, we require continuous clips without camera cuts or scene transitions, clear visuals with identifiable target boundaries, a duration between 10 and 40 seconds excluding static images, and at least one eligible target object. For target objects, we focus on humans, animals, and their parts whose boundaries are distinguishable, which appear clearly in the first frame, contain at least one visible and locatable key point for most of the video while allowing brief occlusions or exits, and remain in motion either actively or passively. To emphasize challenging tracking scenarios, each target is further required to satisfy at least one difficulty criterion, including rapid motion, high similarity to other objects, occlusion or temporary disappearance, deformation or significant scale/viewpoint changes, and small object size. For target points, we select center points, corner points, or semantically meaningful keypoints such as eyes, hands, or heads; the keypoint must appear in the first frame and is labeled as occluded when it becomes invisible, while points on spherical objects are placed near the center. We choose keypoints rather than arbitrary points based on two considerations. From the practical application perspective, downstream tasks such as 3D reconstruction and SLAM usually rely on semantically meaningful keypoints, which offer stronger distinguishing and descriptive capability, and a small number of high-quality keypoints is often sufficient. From the annotation cost perspective, arbitrary point annotation is prohibitively expensive in real-world videos. Existing arbitrary-point datasets often rely on indirect annotation sources, such as optical-flow-based trajectory interpolation in RoboTAP or rendering-based point annotations in Kubric and RGB-Stacking. However, these strategies are not directly applicable to our real-world videos from indoor, outdoor, and wild environments, making it difficult to obtain reliable arbitrary-point annotations automatically. Therefore, we ask annotators to select target keypoints and manually label them frame by frame, balancing annotation utility and feasibility.
5.2 Annotation Pipeline
We designed a coarse-to-fine annotation pipeline to obtain high-quality multi-granularity annotations efficiently and consistently. The pipeline progressively converts selected videos into aligned point, box, and mask annotations through four steps. The overall pipeline is demonstrated in Fig. 4.
1) Video Selection. We downloaded a large number of videos from YouTube and instructed annotators to select videos and target objects meeting the above requirements. This step filters unsuitable clips early and keeps clear, trackable, and challenging targets for annotation.
2) Coarse Annotation. Annotators first mark key points and tight bounding boxes on target objects. These annotations provide fine-grained locations and object-level extents, and serve as prompts for generating initial masks.
3) Fine Annotation. To reduce annotator workload, we use SAM [41] to generate rough masks from coarse annotations (points and boxes), so annotators can focus on boundary correction rather than drawing masks from scratch. Then, annotators refine these masks frame by frame with the following requirements:
Figure 4: Annotation pipeline of our Tracking-Any-Granularity dataset. The pipeline first selects qualified videos and targets, then performs coarse point and box annotation, uses SAM to produce coarse masks for annotator refinement, and finally conducts expert completion checks to ensure consistency across point, box, and dense mask annotations.
TABLE III: Comparison of Automatic Annotation in Different Annotation Methods.
(a) Mask automatic annotation
#1 94.4 91.7 97.1 94.6 91.9 97.3
#2 95.0 92.4 97.6 95.0 92.4 97.6
#3 94.5 91.8 97.2 94.7 92.0 97.4
#4 95.0 92.4 97.6 95.0 92.4 97.7
(b) Box automatic annotation
Type + 0th Back. AUCval $P_{Norm}^{val}$ $P^{val}$ AUCtest $P_{Norm}^{test}$ $P^{test}$
#1 84.1 94.1 91.6 84.8 94.8 91.2
#2 83.7 93.9 91.3 84.5 94.7 91.0
#3 83.9 93.8 91.3 84.7 94.5 91.0
#4 83.6 93.6 91.1 84.4 94.4 90.9
(c) Point automatic annotation
Type + 0th Back. Vis. OAval AJval OAtest AJtest
#1 - 89.0 61.7 88.7 61.3
#2 - 89.0 62.3 88.8 61.6
#3 OR 89.3 58.1 89.6 57.3
#4 AND 88.2 62.1 87.7 61.1
#5 OR 89.4 58.8 89.6 57.7
#6 AND 88.4 62.6 87.8 61.4
(d) Mask & Box Visible automatic annotation
@@ 5.2 Annotation Pipeline
#16 VOS 99.14 99.64 99.42 99.52 99.15 99.69 99.37 99.52
Only annotate the visible parts of the present object in each frame. In cases of motion blur, infer the approximate position based on the previous frame to maintain temporal consistency. Masks in adjacent frames should not differ drastically. Ignore transparent or semi-transparent watermarks and subtitles when creating masks; masks can directly cover these elements. Exclude opaque overlays from the mask, such as logos or captions. For containers holding other objects, do not include the contained objects in the mask. The mask should tightly fit the object, neither exceeding nor falling short of its boundaries. Ensure that mask edges are smooth and avoid excessive roughness. Fill in small internal holes, but preserve natural gaps (such as hollowed-out structures) or occlusions caused by other objects. If the initial SAM-generated mask is of very poor quality, annotators may clear it entirely and use color tolerance-based selection to manually annotate the object from scratch.
4) Final Completion. Experts perform a final review to thoroughly assess the accuracy and consistency of all three types of annotations, ensuring that the labeling meets the required standards and that any discrepancies are identified and corrected, particularly for challenging scenarios such as occlusions and motion blur.
TABLE IV: Comparison of our datasets with public datasets of three tracking datasets in terms of videos, duration, and annotations.
Video Object Segmentation
@@
BURST [1] 2914 1734 (0.60) 624240 (214) $\geq$480p 6 600157 (206) × × Semi-Automatic Multi-Task
LVOS [31] 220 351 (1.60) 126280 (574) 720p 6 156432 (711) × × Manual Long-term
LVOS v2 [32] 720 823 (1.14) 296401 (412) 720p 6 407945 (567) × × Manual Large-scale, long-term
MOSE [21] 2149 443.62 (0.21) $\sim$159600 (73) 1080p 6 431725 (201) × × Semi-Automatic Complext scenarios
YoutubeVOS-19 [73] 4453 334.8 (0.08) 120532 (27) 720p 6 197272 (44) × × Manual Large-scale
VOST [60] 713 252 (0.35) 75547 (106) 1080p 5 175913 (247) × × Semi-Automatic Object transmission
Single Object Tracking
LaSOT [25] 1400 1950 (1.39) 3.52M (2506) 720p 30 × 3.52M × Manual Large-scale, long-term
@@
TrackingNet [51] 30643 8400 (0.27) 14431266 (471) 360p 30 × 14431266 × Semi-Automatic Large-scale
UAV123 [4] 123 62.5 (0.51) 112578 (915) 720p 30 × 112578 × Semi-Automatic Unmanned aerial vehicles
NfS [40] 100 26.58 (0.27) 383K (3830) 720p 240 × 383K × Manual High Frame Rate
OTB-100 [69] 100 32.8 (0.33) 59040 (590) $\geq$360p 30 × 59040 × Manual Real world
TNL2K [63] 2000 691.3 (0.35) 1244340 (622) 720p 30 × 1244340 × Manual Language-based
VastTrack [54] 50610 11664 (0.23) 4.2M (83) 480p-720p 6 × 4.2M × Manual Abundant categories
Point Tracking
@@
PointOdyssey [86] 104 120 (1.15) $\sim$216K (2035) 540p 30 × × 49B (0.471B) Automatic Real world, long-term
TAP-Vid Kinetics [23] 1189 198.17 (0.17) 297250 (250) $\geq$720p 25 × × 4725959 (3974) Semi-Automatic Abitrary point
TAP-Vid DAVIS [23] 30 - (-) 1999 (66.6) 1080p - × × 28824 (960.8) Semi-Automatic Abitrary point
TAP-Vid RGB-Stacking [23] 50 - (-) 12500 (250) 256x256 - × × 303436 (6068.7) Semi-Automatic Abitrary point
Tracking-Any-Granularity 6000 1338.7 (0.22) 2200891 (367) mostly 720p 30 2148716 (358) 2148716 2640987 (440) Semi-Automatic Any Granularity
TABLE V: Statistical analysis of video data from our dataset.
@@
5.3 Data Engine
As shown in Table II, the Tracking-Any-Granularity dataset is annotated across three phases: 1) Phase ①: manual annotation of every frame, totaling 1,000 videos; 2) Phase ②: manual annotation of every 10 frames, totaling 2,000 videos; 3) Phase ③: manual annotation of every 20 frames, totaling 3,000 videos. To increase the size of the dataset while reducing the workload, we adopted a selective annotation strategy in the second and third phases. Instead of manually labeling every video frame, annotators labeled only a subset of frames at varying intervals. After training the model on both public datasets and the fully labeled data from earlier phases, we leveraged the model to automatically annotate the remaining frames. Specifically, each video was divided into multiple clips, with annotators manually labeling the first and last frames of each clip. The annotation of the first frame in each clip served as the initial target state, enabling the model to infer the target state in the intermediate frames.
To further enhance annotation quality, we introduced two optional refinement methods: (1) performing backward tracking and fusing the results with those from forward tracking, and (2) since the target may be absent in some annotated frames, using the first frame of the entire video, which is guaranteed to contain the target, as an additional reference state alongside the first frame of each clip. We evaluated these enhancement methods on the Phase 1 validation and test set to determine the inference setting for each tracking task, as shown in Table III. Specifically, the four subtables represent the evaluation outcomes for the VOS, SOT, PT, and object existence prediction tasks under various settings, respectively. By comparing the results across different settings, we select the configuration highlighted in the gray row as the inference setting for each task.
Although the data engine’s concept is inspired by SAM 2, there are significant differences. The data engine of SAM 2 requires repeated manual correction and re-tracking, which hinders progress. Instead, we decouple manual and automated processes using interval annotations, enhancing efficiency while avoiding error accumulation.
Figure 5: Attribute analysis of video data from our dataset. The figure reports the number of sequences associated with each challenging attribute, covering 18 categories such as motion blur, deformation, occlusion, scale variation, and viewpoint change. Since attributes are not mutually exclusive, one sequence may be assigned to multiple categories.
@@ 5.4 Tracking-Any-Granularity Dataset
Compared with existing datasets, our challenging real-world Tracking-Any-Granularity dataset stands out as the only one providing annotations at all three granularities simultaneously. It comprises 6,000 videos, each annotated with masks, boxes, and points. The majority of the videos have a resolution of $1280\times 720$, with 398 exceptions; their duration ranges from $5.3$ seconds to $110.9$ seconds, and the frame count ranges from $80$ to $3,317$ frames. In total, TAG contains $2.2$ million frames and $1,338.7$ minutes of videos. More detailed statistics are shown in Table V, and annotation examples are shown in Fig. 3. We compare our dataset with numerous public datasets in Table IV, showing our dataset contains significantly more videos and annotations than they do, creating a substantial resource for multi-granularity tracking research.
Scene and Attribute. To enable a more comprehensive analysis of tracking approaches, it is critically important to identify video scenes and attributes of our dataset. Furthermore, we label each sequence with 18 attributes that represent various video challenges, and the detailed definitions and counts are shown in Fig. 5. It is worth noting that these attributes are not mutually exclusive, and a single video may contain multiple challenges. Motion Blur, Deformation, and Partial Occlusion are the most common challenges in our dataset, demonstrating its high level of difficulty.
TABLE VI: State-of-the-art comparison on Video Object Segmentation Task.
[…]
STCN [16] 45.8 45.6 45.9 62.0 58.5 65.4 50.1 46.1 54.1 70.4 65.9 75.0 76.2 72.2 80.2 32.7 22.7 82.8 81.3 85.6 78.3 86.0
AOT-SwinB [79] 54.9 55.3 54.8 73.8 70.1 77.4 60.2 56.2 64.1 78.1 73.1 83.2 80.9 76.4 85.4 39.3 29.4 85.3 84.6 89.5 79.3 87.7
DeAOT-SwinB [80] 58.5 58.0 58.6 72.1 68.5 75.8 61.7 57.6 65.9 79.6 74.8 84.4 81.6 77.3 85.9 43.0 29.4 86.4 85.4 90.3 80.6 89.3
XMem [15] 52.3 51.9 52.3 64.7 61.8 67.5 56.3 52.1 60.5 74.4 70.1 78.6 75.7 71.8 79.6 37.9 25.0 85.5 84.3 88.6 80.5 88.7
DEVA [14] 56.1 55.4 56.2 72.2 68.6 75.9 66.4 62.2 70.6 77.9 73.1 82.6 82.1 78.0 86.1 40.9 27.3 86.3 85.3 89.8 80.6 89.2
Cutie-base+ [13] 55.8 55.6 55.8 71.4 68.6 74.3 66.2 62.3 70.1 79.0 75.0 83.0 83.8 80.0 87.7 44.7 32.7 86.9 86.2 90.7 81.6 89.2
[…]
JointFormer [84] - - - 71.7 68.8 74.7 69.7 65.8 73.6 76.6 72.8 80.5 79.1 75.5 82.7 - - 87.0 86.1 90.6 82.0 89.5
Ours 66.4 66.5 66.4 82.2 78.7 85.7 74.6 70.6 78.6 87.4 84.2 90.7 87.9 84.9 90.9 45.2 25.6 87.1 85.8 82.5 90.0 90.3
TABLE VII: State-of-the-art comparison on Single Object Tracking Task.
Methods GOT10ktest $_{SOT}^{test} TAG$$_{SOT}^{val}$ TrackingNet VastTrack
AO $SR_{0.5}$ $SR_{0.75}$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$
[…]
SAMURAI [78] 79.6 90.8 72.9 - - - - - - 85.3 - - - - -
[…]
Ours 80.7 89.7 77.8 78.0 85.7 81.5 78.2 86.2 82.0 86.0 90.1 87.3 55.0 65.5 60.4
TABLE VIII: State-of-the-art comparison on Point Tracking Task. model and model means original offline trackers and modified online trackers.
Method Key Point Arbitrary Point Method Key Point Arbitrary Point
BADJA Perc.val $_{PT}^{test} TAG$$_{PT}^{val}$ DAVIS RGB BADJA Perc.val $_{PT}^{test} TAG$$_{PT}^{val}$ DAVIS RGB
pips [29] 64.2 41.5 19.0 19.8 40.9 28.5 TAPTR [44] 69.1 59.4 23.7 23.8 61.2 58.0
pips [29] 56.9 27.2 12.3 12.2 26.3 20.8 TAPTR [44] 63.0 48.9 20.4 19.0 52.3 39.4
pips++ [86] 56.6 62.6 20.9 23.1 59.8 58.5 TAPIR [24] 64.2 59.6 21.3 24.6 56.8 50.4
CoTracker [38] 65.3 59.8 23.3 22.3 60.9 63.1 LocoTrack [17] 68.7 67.1 25.2 30.2 62.7 69.2
CoTracker [38] 55.1 48.9 18.8 18.1 50.8 46.1 Track-On [2] 69.7 69.7 24.8 25.8 64.5 64.5
CoTracker3 [37] 72.7 71.3 29.6 29.1 65.6 70.6 Ours 72.9 66.2 35.3 37.7 56.1 59.0
CoTracker3 [37] 66.3 66.3 25.8 24.9 59.2 63.6
Dataset Splits. As summarized in Table II, Phase ① contains 1,000 fully annotated videos with dense manual labels on every frame, making it the most reliable source for evaluation splits. From this phase, we selected 150 validation videos and 150 test videos with stratified sampling based on both category and source, ensuring a balanced distribution and reducing potential bias for fair evaluation across diverse video types.
6 Experiments
Figure 6: Comparison between our model and various SOTA methods on video tracking benchmarks at three granularities. The three groups show representative results for mask-level video object segmentation, box-level single object tracking, and point-level online point tracking. Across scenarios with deformation, occlusion, fast motion, and appearance changes, SAM 2++ produces more stable target states and accurate localization than task-specific baselines, preserving object boundaries, bounding boxes, and semantic points across frames.
Figure 7: Examples from SAM 2++ results on video benchmarks at various granularities. The same model outputs segmentation masks, bounding boxes, and target points from different prompts, while maintaining consistent predictions across diverse categories and motion patterns.
6.1 Comparison to state-of-the-art on three tasks
Video Object Segmentation. The comparisons between our model and previous semi-supervised VOS methods are demonstrated in Table VI, including YoutubeVOS-19 [73], MOSE [21], LVOS-v2 [32], BURST [1], VOST [60], VISOR [20], and our TAG dataset. For evaluation, we use standard metrics [55] in most benchmarks: Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$. In the YouTubeVOS benchmark, $\mathcal{J}$ and $\mathcal{F}$ are computed for ”seen” and ”unseen” categories separately, and $\mathcal{G}$ is the averaged $\mathcal{J\&F}$ for both seen and unseen classes. In LVOS, the standard deviation $\mathcal{V}$ of the average score of $\mathcal{J}$ and $\mathcal{F}$ assesses temporal stability, while VOST additionally reports $\mathcal{J}_{tr}$ for the last 25% frames to evaluate robustness after object transformation. The BURST benchmark is evaluated with Higher Order Tracking Accuracy (HOTA) [50], which balances frame-level detection and temporal association accuracy. Results show that our model outperforms individual video object segmentation models.
Single Object Tracking. We compare the performance of our proposed model on three benchmarks in Table VII, including TrackingNet [51], GOT-10k [34], TNL2K [63], VastTrack [54] and our TAG dataset, and all compared models are trained on four datasets. For most benchmarks, we evaluate performance with Area Under the Curve (AUC), normalized precision (${P}_{Norm}$), and precision (P{P}), which measure the average accuracy of center, size, and scale between predicted and ground-truth bounding boxes. For the GOT-10k benchmark, we choose Average Overlap (AO) and success rate (SR) as indicators, where AO denotes the average overlap between all ground-truth and estimated boxes, and SR measures the percentage of successfully tracked frames whose overlap exceeds a threshold (e.g., 0.5). Experiments demonstrate that our model consistently outperforms existing methods across nearly all benchmarks. Crucially, unlike specialized SOT trackers that focus on a single task and employ a pre-cropping process, our framework is an end-to-end, unified tracker.
Online Point Tracking. We compare our method to prior works in Table VIII on four benchmarks, including BADJA [5] and Perception Test [52] for key point tracking, TAP-Vid [23] for arbitrary point tracking, and our TAG dataset. For TAP-Vid [23], Perception Test [52], and our TAG dataset, we report Occlusion Accuracy (OA) and Average Jaccard (AJ). For the BADJA benchmark, we adopt the Percentage of Correct Keypoint-Transfer (PCK-T). We measure these benchmarks in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries. However, most of the current methods are offline trackers, which process long-temporal window frames or even the entire video to be able to see the future frame, and do not match the online setup required by real applications. For a fair comparison, we modified their input so that there is no future information inside the window, denoted as model, to enable inference online. Experimental results show a substantial decrease in model performance when the input data is switched from offline to online. The comparative analysis reveals the effectiveness of our approach on keypoint tracking benchmarks, which surpasses competing models. Furthermore, although our model is trained on keypoint datasets, it demonstrates generalization capability on arbitrary point tracking datasets.
Comparison with TAG model. TAG [30] shares our goal of unified mask, box, and point tracking, so we compare the two approaches to clarify our contributions. First, TAG is an offline model that processes multiple frames as a clip, which differs from the mainstream online frame-by-frame setting, may leak future information, and is mainly applicable to pre-recorded videos. In contrast, our model uses only past information for the current frame, making it suitable for video streams. Second, TAG converts point coordinates into a {0, 1} mask, providing limited target information. Our method combines point coordinates with a (0, 1) Gaussian mask, using precise coordinates for localization and the Gaussian mask to highlight the target in a form consistent with MaskDecoder outputs and MemoryEncoder inputs. For box tracking, TAG converts a box into a square mask, which may confuse the target region with the background. Third, TAG is trained only on public datasets, while our Data Engine supports both model training and annotation expansion, yielding a large-scale dataset with three granularities and a stronger unified model. Most importantly, TAG focuses on jointly encoding targets of different granularities in an offline clip-level setting. When moving to the next clip, it still relies on prompts in their original form and must use the last-frame prediction from the previous clip as the next prompt, even when that prediction is unreliable, which can cause error accumulation and makes temporary disappearance difficult to handle. Our online formulation instead requires converting predictions into memory representations for subsequent frames. Memory provides richer target features than raw prompts and, with selective capability and memory diversity, improves tracking stability. To support this across granularities, we introduce Task-Adaptive Memory to unify predictions of different forms into memory representations. Thus, the difference from TAG is not only about task coverage, but also about motivation and core design. As shown in Table IX, our model significantly outperforms TAG on all three tasks.
TABLE IX: Comparison with TAG Models.
Methods VOST LaTOT TOTB CroHD
TAG [30] 31.3 35.3 74.4 57.1
Ours 45.2 37.4 83.8 66.6
TABLE X: Comparison with Unified Models.